Find the mean of the probability distribution of the random variable X, which can take only the values 3,5, and 7, given that P(3)= 7/30, P(5)=1/3, and P(3)= 13/30.

1
Expert's answer
2022-03-09T08:07:16-0500

"\mu=3\cdot\cfrac{7}{30}+5\cdot\cfrac{1 }{3}+7\cdot\cfrac{13} {30} ="


"=\cfrac{3\cdot7+5\cdot10+7\cdot13} {30} ="


"=\cfrac{162} {30} =5.4."